# Total trades, intersection matrices and Specht modules

### v1

@@ 4.2. Weyl modules and Pieri’s rule
Since the characteristic of $\mathbb{K}$ is zero, $K_{\mu}$ is an irreducible $G$-module. Moreover, if $\lambda$ and $\mu$ are distinct partitions, then the irreducible modules $K_{\lambda}$ and $K_{\mu}$ ar not isomorphic.
@@ 4.3. Straightening row semistandard tableaux
Given $Z:={1}^{(a_{1})}{2}^{(a_{2})}\cdots{N}^{(a_{N})}\otimes{1}^{(b_{1})}{2}^{(b_{2})}\cdots{N}^{(b_{N})}\in D_{\mu_{1}}\otimes D_{\mu_{2}}$, let $T(Z)$ be the Young tableau with first row consisting (from left to right) of $a_{1}$ $1$’s, …, $a_{N}$ $N$’s and second row consisting (from left to right) of $b_{1}$ $1$’s, …, $b_{N}$ $N$’s. We will need the following theorem.
@@ 5. Images of intersection matrices
Using the definitions of $\Psi_{l}$ and $\pi_{j}$ (for the later see (4.2)) we have
@@
$A=\bigoplus_{j\in J(t,k,n)}S^{(n-j,j)},$
where $J(t,k,n)$ is as in Theorem 5.3. In particular, the rank of the map $A$ is equal to
@@
(1) Corollary 3.6 generalizes [6, Corollary 14] that proves an expression for the rank of the individual intersection matrices $U_{t,k,l}$ by completely different methods. (2) As a very special case, Corollary 3.6 contains the statement that the inclusion matrices $W_{t,k}=U_{t,k,t}$ have maximal rank and hence we have another proof of this classical result [9, 1, 3, 21].

### v2

@@ 4.2. Weyl modules and Pieri’s rule
Since the characteristic of $\mathbb{K}$ is zero, $K_{\mu}$ is an irreducible $G$-module. Moreover, if $\lambda$ and $\mu$ are distinct partitions, then the irreducible modules $K_{\lambda}$ and $K_{\mu}$ are not isomorphic.
@@ 4.3. Straightening row semistandard tableaux
Given $Z:={1}^{(a_{1})}{2}^{(a_{2})}\cdots{N}^{(a_{N})}\otimes{1}^{(b_{1})}{2}^{(b_{2})}\cdots{N}^{(b_{N})}\in D_{\mu_{1}}\otimes D_{\mu_{2}}$, let $T(Z)$ be the Young tableau with first row consisting (from left to right) of $a_{1}$ 11’s, …, $a_{N}$ NN’s and second row consisting (from left to right) of $b_{1}$ 11’s, …, $b_{N}$ NN’s. We will need the following theorem.
@@ 5. Images of intersection matrices
Using the definitions of $\Psi_{l}$ and $\pi_{j}$ (for the latter see (4.2)) we have
@@
$\bigoplus_{j\in J(t,k,n)}S^{(n-j,j)},$
where $J(t,k,n)$ is as in Theorem 5.3. In particular, the rank of $A$ is equal to
@@
(1) Corollary 5.6 generalizes [6, Corollary 14] that proves an expression for the rank of the individual intersection matrices $U_{t,k,l}$ by completely different methods. (2) In Corollary 5.6 consider $l=t$, $c_{0}=\cdots=c_{t-1}=0$ and $c_{t}=1$. Then $A=U_{t,k,t}=W_{t,k}$ is the inclusion matrix. Let $j\in\{0,1,\dots,t\}$. Then from Definition 5.1 we have
$\lambda_{j}(t,k,n;t)=\sum_{s=0}^{j}(-1)^{j-s}\tbinom{j}{s}\tbinom{k-s}{t-s}\tbinom{n-k-j+s}{-j+s}=\tbinom{k-j}{t-j}$,
where in the last equality we used our convention on binomial coefficients that $\tbinom{n-k-j+s}{-j+s}=0$ if j+s<0-j+s<0. Thus with the notation of Corollary 5.6 we obtain $J(t,k,n)=\{0,1,\dots,t\}$ and the rank of the matrix $A=W_{t,k}$ is equal to
$\sum_{j\in J(t,k,n)}\big(\tbinom{n}{j}-\tbinom{n}{j-1}\big)=\big(\tbinom{n}{0}-\tbinom{n}{-1}\big)+\big(\tbinom{n}{1}-\tbinom{n}{0}\big)+\cdots+\big(\tbinom{n}{t}-\tbinom{n}{t-1}\big)=\tbinom{n}{t}$.
In other words we obtain that the inclusion matrix $W_{t,k}$ has maximal rank. This is a well known result [9, 21, 1, 3].
6. Acknowledgments
We are grateful to the referee for carefully reading the manuscript and for providing useful suggestions.